OpenStudy (anonymous):

how do you solve for x-intercept in: y = (x - 5)^2 - 4?

OpenStudy (anonymous):

x-intercept => y=0 so (x-5)^2 - 4 = 0 => (x-5)^2 = 4 => x-5 = +- 2 => x= 7 and x= 3

OpenStudy (anonymous):

wait.. what'd you do with the square?

OpenStudy (anonymous):

sqrt 2 sides of equation sqrt (x-5)^2 = sqrt 4 (x-5)=+-2